A research scientist wants to know how many times per hour a certain strand of bacteria reproduces. The mean is found to be 5.25.2 reproductions and the population standard deviation is known to be 2.12.1. If a sample of 775775 was used for the study, construct the 90%90% confidence interval for the true mean number of reproductions per hour for the bacteria. Round your answers to one decimal place.
Answer #1

Solution :

Given that,

Sample size = n = 775

Z\alpha/2 = 1.645

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.645 * (2.1 / \sqrt 775)

= 0.1

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

5.2 - 0.1 < \mu < 5.2 + 0.1

5.1 < \mu < 5.3

(5.1 , 5.3)